average wavelength of light emitted by a 100 watt bulb is 5,000  A^0 . the number of  emitted photons per second :
(1) 5*10^17   (2) 2.5*10^22   (3) 3*10^23   (4) 2.5*10^19

Grade:12

1 Answers

Venkat
5 years ago
Power of bulb is given as 100 watt. Calculate the energy first and then no.of photons.
Energy of one photon is
 
E=h\frac{c}{\lambda } = \frac{6.626*10^{-34}*3*10^{8}}{5000*10^{-10}}=3.975*10^{-18}
No.of photons emitted per second is
=\frac{100}{3.97856*10^{-18}} = 2.515*10^{19}